3. La Besurta

’IBONS DE GORGUTES’ (GORGUTES LITTLE MOUNTAIN LAKES) 1 hour and 30 min. respectively

From the end of the main road over the Pllan de l’Ospital, sets off a track that ascends over the left slope of the valley to the North. After 30 minutes it crosses the Gorgutes ravine, creating a wide zig-zag on the opposite slope. After a 1 hour walk, you reach a spot over the little ibones of the Solana de Gorgutes on your right under the wall of La Tuca de la Montañeta. If you follow the track on your left you will get to the Ibón de Gorgutes (2,313 m. high). You will enjoy some stunning views from this point.

If you circumvent the ibón on its right, you will reach the Puerto de la Glera (2,364 m. high) in about 15 minutes. This is a border port over the town of Luchón. If you go up the left slope you will contemplate and admire the whole Luchón valley and the ski resort of Superbaquères.